OpenStudy (anonymous):

solve for x 2(x-4)+(x+2)=(x+2)(x-4)

OpenStudy (anonymous):

\[2x-8+x+2=x^2-2x-8\] \[3x+6=x^2-2x-8\] \[x^2-5x-14=0\] \[(x-7)(x+2)=0\] \[x-7=0\] \[x=7\] \[x+2=0\] \[x=-2\]

OpenStudy (anonymous):

that is not correct?

OpenStudy (anonymous):

ok let me try it again. maybe an algebra mistake

OpenStudy (amistre64):

tsk tsk tsk.....

OpenStudy (anonymous):

sorry master...

OpenStudy (anonymous):

oh damn it was -6 not 6!

OpenStudy (amistre64):

2(x-4)+(x+2)=(x+2)(x-4) 2x-8 +x+2 = x^2 -2x -8 3x -6 = x^2 .......

OpenStudy (amistre64):

how many times I gotta clean up your messes lol

OpenStudy (anonymous):

\[3x-6=x^2-2x-8\] \[x^2-5x-14=0\]

OpenStudy (anonymous):

another mistake!

OpenStudy (anonymous):

\[x^2-5x-2=0\] whew

OpenStudy (amistre64):

the cape and cowl are going back to the store ......

OpenStudy (anonymous):

answer is \[x=\frac{5\pm\sqrt{33}}{2}\]

OpenStudy (anonymous):

damned algebra kills me every time. where did i get a +6 from??

OpenStudy (anonymous):

now pt jessica it is right.

OpenStudy (amistre64):

trying to type, latex, and math at the same time; you lack discipline my child...

OpenStudy (anonymous):

trust me or the master puts me back in the dungeon

OpenStudy (anonymous):

i am going to get you to use latex eventually...

OpenStudy (amistre64):

as long as my extra finger aint doin nuthin it might as well take up the latex lol

OpenStudy (anonymous):

so, what is the answer again?

OpenStudy (anonymous):

\[2(x - 4) + (x+2) = (x+2)(x-4)\] \[2x-8 + x + 2 = x ^{2} - 4x + 2x - 8\] \[x ^{2} - 5x -2 = 0\] \[\Delta=25-4*1*(-2) = 25 + 8 = 33\] \[x=-(-5) + \sqrt{33}/2 or x=-(-5)-\sqrt{33}q2\]

OpenStudy (anonymous):

\[solution 1, x=(\sqrt{33}+5)/2 solution 2, x=(5-\sqrt{33})/2\]